OpenStudy (2yoututu11):

Each side of an equilateral triangle is 32 centimeters, as shown. What is the approximate height h of the triangle? A. 16 cm B. 28 cm C. 30 cm D. 32 cm

OpenStudy (anonymous):

Use the Pythagorean theorem with\[\left(\frac{32}{2}\right)^2+h^2=32^2.\]Solve for \(h\).

OpenStudy (anonymous):

the height of that equilateral triangle is basically the longer leg of the 30-60-90 triangle. using the properties of the 30-60-90 triangle, the side you want is just (the shortest side)*sqrt3